OpenStudy (anonymous):

Consider the following system of equations. y = 4x - 4 y - 4x = 3 What can you conclude about the system of equations?

OpenStudy (anonymous):

Consider the equation of a straight line: y=mx+b m = the slope. if two lines have the same slope they are parallel lines.

OpenStudy (anonymous):

I believe that these system of equations: y=0 so the next equation would read: 0-4x=3

OpenStudy (anonymous):

rewrite your second equation so that the x term is on the otherside and see if the slope is the same.

OpenStudy (anonymous):

x=-3/4

OpenStudy (anonymous):

so the systems are on the same line, right? So wouldn't that make them dependent?

OpenStudy (anonymous):

They are not on the same line as they cross the y axis at different points, -4 and 3.

OpenStudy (anonymous):

Parallel lines are inconsistent.

OpenStudy (anonymous):

I see! Thanks!

OpenStudy (anonymous):

:)